What is 15-(4m-5)=32?
Answer by LilSkittleMd(119)

15-(4m-5)=32
First distribute what is in the ( )
15-4m+5=32
20-4m=32
Subtract 20 from both sides
-4m=12
Divide both sides by -4
m=-3